Assume the random variable X is normally distributed with mean as 41 and the standard deviation as 7. Find the 5th percentile.
Answer #1

P(X < x) = .0.05

or, P((X - )/ < (x - )/) = 0.05

or, P(Z < (x - 41)/7) = 0.05

or, (x - 41)/7 = -1.645

or, x = -1.645 * 7 + 41

or, x = 29.485